OpenStudy (anonymous):

f(x) = 3x + 2; g(x) = 3x - 5 Find f/g.

jhonyy9 (jhonyy9):

so f/g =(3x+2)/(3x-5) yes ?

jhonyy9 (jhonyy9):

so can you calcule it ?

OpenStudy (anonymous):

yeah but i have choicess. i got -2/5 but thats not right

jhonyy9 (jhonyy9):

3x+2 / 3x -5 = 1 3x-5 ---- 0 7 (3x+2)/(3x-5)=1 and remander 7

OpenStudy (anonymous):

yeah, thats not an answer.

OpenStudy (anonymous):

i attached the answers..

hartnn (hartnn):

denominator can't equal 0 so 3x-5 \(\ne\)0 sp x\(\ne\) ?

hartnn (hartnn):

*so

OpenStudy (anonymous):

...?

jhonyy9 (jhonyy9):

what value not can getting x ?

OpenStudy (anonymous):

the answers are right there..

hartnn (hartnn):

3x-5\(\ne\)0 3x\(\ne\)5 x\(\ne\)5/3

OpenStudy (anonymous):

i dont understand what to do?

OpenStudy (anonymous):

thank you